A 56-kg rider on a slingshot ride gets launched from the rest to a final velocity of 13 m/s in a time of 1.4 seconds. What is the force that acts on the rider?
1
Expert's answer
2019-06-10T09:47:43-0400

Solution. Change in momentum is proportional to the net external force and the time over which a net force acts. Hence


"\Delta p = F\Delta t ."

Therefore


"F=\frac {\Delta p} {\Delta t}."

From the definition of the linear momentum get


"\Delta p = m(v_2 - v_1)."

As result


"F=\frac { m(v_2 - v_1)} {\Delta t}."

According to the condition of the problem m=56kg, v2=13 m/s, v1=0 m/s, change in time t=1.4s. Hence


"F=\frac { 56(13 -0)} {1.4}=520N."

Answer. 520 N.